solve for x
2 square root of x over c = y

Answer by RAY100(1637)

2sqrtx/c=y
.
cross multiply
.
2sqrtx = cy,,,,,,,divide by 2
.
sqrtx = cy/2,,,, square both sides
.
x= c^2y^2/4
.
check,,,{2 sqrt c^2 y^2/4}/c = y
.
2(cy/2c)=y,,,,,ok
.